Martha works for a prominent technology company. Her company just paid a $1.50 dividend per share. The required return for her company’s stock is 12%.

Question:

Consider the following information. Suppose Martha’s company is expected to increase dividends by 12% in one year, and by 8% in two years. After that, her company’s dividends will increase at a rate of 6% indefinitely. If the last dividend was $1.50 and the required rate of return in 12%, what is the current price of the stock?
Answer #1

D1=(1.5*1.12)=1.68

D2=(1.68*1.08)=1.8144

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(1.8144*1.06)/(0.12-0.06)

=32.0544

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.68/1.12+1.8144/1.12^2+32.0544/1.12^2

=$28.5